OpenStudy (anonymous):

Find the volume of a cube that has a total area of 96 sq. in. V = cu. in.

OpenStudy (anonymous):

6 sides, let length of one of them be x in. Area of 1 side, x by x = x^2 sq in. 6 sides, 6x^2 = 96 so x^2 is 16 and x = 4 inches. Volume 4*4*4 = 4^3 = 64 Cubic in.